I recently sent you a substitution problem , but it actually a addition method problem.Sorry. The problem was:
x-4y=5
x-2y=1

My answer was -3, -2. Is that correct?
Thankyou

Answer by checkley71(8403)

X-4Y=5
-X+2Y=-1
-----------------------------
-2Y=4
Y=4/-2
Y=-2 ANSWER.
X-4*-2=5
X+8=5
X=5-8
X=-3 (-3,-2) IS THE CORRECT ANSWER.
-(-3)+2(-2)=-1
3-4=-1
-1=-1